# E-scooter effects on public transport demand: a case study in Santiago, Chile

### v1

@@ 6 Discussion
The results of this study provide a comprehensive view of the impact of e-scooters on public transportation in Santiago, Chile, considering both generations and attraction scenarios. By unifying these findings, three key aspects emerge: the regional variability of e-scooter impacts, the complementary and competitive role of e-scooters, and strategic implications for urban mobility planning.
Regional Differences
In both scenarios, we see that e-scooters affect public transportation differently in each region, based on the percentage changes per hour and per zone. In the Central Region, e-scooters compete with public transport overall, reducing combined bus and metro trips in both the generation (-21.38% per hour per zone) and attraction (-72.06% per hour per zone) scenarios. However, this region also shows a substantial increase in bus trips associated with e-scooter usage, with increases of 76.39% in trip generation and 38.12% in trip attraction, both measured per hour per zone. This pattern suggests that while e-scooters may displace some metro trips, they are facilitating access to and from buses, particularly in areas where walking distances might be longer or less convenient.
In contrast, the Intermediate Regions show predominantly complementary effects of e-scooters for both transport modes, metro and buses. In the Intermediate Region, e-scooters increase combined bus and metro usage by 56.83% in trips generation and 63.85% in trips attraction, both per hour and per zone, with a slight preference for metro usage (70.05% in generation and 62.17% in attraction). In the Peripheral Region, scooters significantly increase the generation of public transport trips, as well as the overall generation and attraction of bus trips. The impact is noticeable, with a 40.17% increase in generated trips, and in the case of bus trips, the increases are 84.64% in generation and 77.29% in attraction, all per hour and per zone. These results highlight the role of e-scooters in areas with varying levels of public transport coverage, where they can complement existing services by improving accessibility and connectivity.
By associating public transport generation trips with first-mile trips made on scooters, it is observed that scooters serve as a first-mile mode for accessing both buses and the metro in the Intermediate region, while facilitating only bus access in the Peripheral and Central regions. When considering that public transport attraction trips can also be linked to last-mile scooter trips, it becomes evident that they are used as a last-mile mode for exiting buses and the metro in the Intermediate region, and once again, only for buses in the Peripheral and Central regions.
These findings are consistent with those of Luo et al. (2021), who also found that e-scooters can either compete with or complement public transportation depending on the urban context and the availability of other transport options. However, our study goes further by considering not only buses but also the metro, providing a more comprehensive understanding of how e-scooters can be integrated into multimodal transport systems across different urban regions.
Differential Impacts on Transport Modes
An intriguing pattern emerged in our results, particularly in the Central Region. While e-scooters were associated with a decrease in combined bus and metro usage, they showed a positive impact on bus trips alone. This seemingly contradictory finding warrants further exploration. One possible explanation is that e-scooters are primarily replacing short-distance metro trips, while simultaneously improving access to bus stops for longer journeys. Additionally, the flexibility of e-scooters might allow users to avoid peak crowding times on metro services, a factor known to influence public transport choice and user experience (Tirachini et al., 2013). This could lead to an overall reduction in combined public transport trips, but an increase in bus usage specifically, as buses might be perceived as less crowded or more accessible when combined with e-scooter use.
Complementary and Competitive Role of E-Scooters
E-scooters demonstrate a dual role, acting as both complements and competitors to public transportation depending on the context. In the Central Region, with its high density and well-developed transport infrastructure, e-scooters seem to compete more directly with public transport. Specifically, e-scooters in the Central Region are associated with a reduction in combined bus and metro trips in both the generation (-21.38% per hour per zone) and attraction (-72.06% per hour per zone) scenarios. However, they also contribute to a substantial increase in bus trips, with rises of 76.39% in generation and 38.12% in attraction trips, both per hour per zone. This suggests that while e-scooters may replace public transport for certain trips, they also provide significant connections for bus users.
In contrast, in the Intermediate and Peripheral Regions, e-scooters clearly play a complementary role by connecting users to public transport and extending the reach of the urban transport system. Given the absence of metro stations in the Peripheral Regions, the results indicate that scooters effectively bridge gaps in transport accessibility, especially for bus trips. In the Intermediate Region, this pattern aligns with studies like Weschke (2023) in Berlin, which found a notable increase in e-scooter demand near metro stations, highlighting the complementary role scooters can play in connecting users to other modes of public transport.
The Peripheral Region in particular benefits from improved bus access, where the increase in public transport connectivity is crucial due to the limited transport options compared to the Central Region. This observation supports the idea that e-scooters can serve as last-mile solutions in areas with less developed transport infrastructure. In the Intermediate Region, a similar complementary pattern is observed, reinforcing e-scooters’ role in facilitating access to public transport.
While our findings highlight the complementary nature of e-scooters in both Intermediate and Peripheral Regions, they differ somewhat from studies in U.S. cities like Louisville and Nashville Ziedan et al. (2021a, b), which reported that e-scooters sometimes reduced bus demand, particularly for routine trips. However, as in our findings, these studies also acknowledged that scooters could increase bus demand for specific routes or social trips. This suggests that the relationship between e-scooters and public transport is highly context-dependent, varying by city structure, transport infrastructure, and user behavior.
@@
Investing in and promoting sustainable travel modes such as cycling and public transport is one of the most promising strategies for reducing transport-related carbon emissions (see e.g. Brand et al., 2021; Amelung et al., 2019; de Nazelle et al., 2010; Keall et al., 2018). Our study contributes to this goal by providing evidence on the effectiveness implementing e-scooters to promote the use of one of these sustainable travel modes, public transport. Specifically, the results of this study have important implications for urban mobility planning in Santiago and other cities with similar characteristics. It shows that the evaluation of e-scooter deployment should be done separately for different areas of a city. In this study we show that in peripheral and moderately dense areas, e-scooters offer an effective solution for improving access to public transport, reducing car dependency, and fostering a more integrated and efficient transport system. This study suggests to policy-makers to prioritize first the installation of e-scooters in zones with less level of accessibility to public transportation in order to incentives the use of it. In this context, policies that promote the integration of e-scooters with public transport, such as the installation of scooter stations near bus stops and metro stations, could maximize the observed benefits.
Our analysis highlights the value of using track-and-trace data (Harrison et al., 2020) from smart cards and GPS data from public transport and e-scooters to understand and influence travel behaviour. This approach enables policy makers to develop targeted, data-driven strategies to promote sustainable travel modes. Coupled with the observed variability in the impact of e-scooters across different travel contexts (different areas of the city and different travel modes), passive transport data allows the identification of specific travel contexts where the effectiveness of interventions such as e-scooter deployment is particularly high compared to other travel contexts. In this context, our study suggests that it is highly recommended to develop transport policies aimed at defining and implementing e-scooter systems targeted at commuters living in peripheral or intermediate areas.
For these strategies to be effective, it is crucial that urban planners and transport authorities have access not only to data from public transport systems but also to data from emerging modes of transport such as e-scooters. Understanding the patterns of e-scooter usage, their interaction with public transport, and the specific needs of different urban areas requires comprehensive data collection and analysis. The integration of such data into urban planning processes can help in designing more effective and responsive transport systems that cater to the diverse needs of the population, enhance multimodal connectivity, and promote sustainable urban mobility.
Limitations
Our study, while offering significant contributions to understanding the interaction between e-scooters and public transportation in Santiago, presents some limitations. First, our focus on a single urban context limits the generalizability of the results to other Latin American cities with different geographical, socioeconomic, and infrastructural characteristics. Additionally, we did not consider the impact of external factors such as seasonal changes, or extraordinary events that could influence the use of public transportation and e-scooters.
Although the use of the difference-in-differences (diff-in-diff) method improves our ability to infer causality, this approach can still be limited by unobserved factors that vary over time and could affect our results, such as changes in the quality of public transport service, fluctuations in the local economy, or variations in user safety perceptions. Additionally, our study calculates short-term effects and recognizes that the long-term complementarity of public transportation and micromobility might be underestimated as user behavior adjusts over time. Assessing their environmental impact and long-term viability as a mobility option is crucial. Future studies could benefit from additional experimental or longitudinal designs to address these limitations and provide a more robust understanding of the observed dynamics.
@@ 7 Conclusions
This study methodically explores how shared e-scooters influence public transport usage in Santiago, Chile, by analyzing trip origins and destinations. The findings indicate that e-scooters can both complement and compete with public transport, depending on the urban context and specific transport modes involved.
7.1 Summary of Key Findings
  • 1.
    Regional Variability: The impact of e-scooters on public transport differs significantly across regions. In the Central Region, e-scooters tend to compete with the metro system while complementing bus services, suggesting their role in improving access to buses in densely populated areas. In contrast, in the Intermediate and Peripheral Regions, e-scooters primarily complement public transport, enhancing accessibility and extending the reach of the transport system, particularly in areas with less developed infrastructure.
  • 2.
    Dual Roles: E-scooters serve both complementary and competitive roles in urban mobility. Their ability to either enhance or compete with public transport modes highlights the need for region-specific policies that account for these dynamics.
7.2 Strategic Implications
This investigation underscores the importance of integrating e-scooter usage data into urban mobility planning. By incorporating comprehensive data from various modes of transportation, urban planners and transport authorities can design more efficient systems that enhance multimodal connectivity and address diverse population needs.
7.3 Implications for Policy and Practice
  • 1.
    Data Integration: Urban mobility strategies should include data from both traditional public transport systems and new transportation modes like e-scooters. This integration is essential for understanding usage patterns, optimizing transport networks, and fostering sustainable urban mobility.
  • 2.
    Tailored Strategies: In Santiago de Chile, region-specific strategies are needed to manage the role of e-scooters in urban transport. In central areas, e-scooter policies can focus on enhancing their role alongside bus services. In intermediate regions, fostering synergy between e-scooters, metro, and bus networks can improve transport efficiency. In peripheral regions, promoting e-scooter use, particularly in combination with bus services, can enhance transport accessibility. For other cities, it is important to study the specific configuration and distribution of their transport systems to adapt these strategies to local contexts.
By addressing these strategic areas, cities like Santiago de Chile, as well as others, can develop more nuanced and effective urban mobility frameworks that balance the roles of e-scooters and public transport.
Acknowledgments
We thank Gonzalo Cortés from Scoot Networks for facilitating the data for this study. D. Opitz and M. Rivas were partially funded by ANID Fondecyt de Iniciación #11220799. E. Graells-Garrido was partially funded by National Center for Artificial Intelligence CENIA FB210017, Basal ANID. N. Meza was partially funded by FCFM starter project “Modelo Eficiente de Fusión de Datos Urbanos.”
@@ Appendix C Results of Negative Binomial Models for Evaluating Generation of Public Transport Trips
Morning peak 0.5638 (41.678) *** 0.2558 (18.403) *** -0.0793 (-2.284) *
Lunch -0.1860 (-12.482) *** -0.0128 (-0.843) 0.1897 (5.074) ***
Afternoon peak -0.1729 (-13.114) *** 0.1939 (14.360) *** 0.5787 (16.969) ***
Afternoon valley -0.1559 (-11.882) *** 0.1558 (11.573) *** 0.4729 (13.886) ***
Night -1.7074 (-118.555) *** -2.3861 (-169.260) *** -2.7928 (-86.563) ***
Night valley -1.1541 (-81.095) *** -0.5905 (-41.059) *** -0.3549 (-10.062) ***
@@
Friday 0.0435 (4.034) *** 0.0449 (4.096) *** 0.0446 (1.669)
Saturday -0.4696 (-42.194) *** -0.4781 (-42.562) *** -0.5387 (-19.821) ***
Sunday -0.8738 (-75.446) *** -0.9108 (-78.426) *** -1.0786 (-38.667) ***
Treatment period
Year 2019 0.0166 (2.231) * -0.0470 (-6.107) *** 0.0200 (1.087)
Treatment zones
Scooter trips destination 0.4079 (3.634) *** 0.5880 (25.502) *** -0.2258 (-2.867) **
@@
Log(working population) 0.1564 (35.042) *** 0.2813 (47.675) *** 0.3028 (18.138) ***
Relative educational level 0.1692 (34.307) *** 0.2125 (44.320) *** 0.5689 (44.883) ***
@@
Table 8: Results of negative binomial models for evaluating generation of bus and metro trips
@@
Morning peak 62.753 (1.580) *** 43.412 (2.378) *** -47.985 (21.022) **
Lunch -20.707 (1.665) *** -2.164 (2.568) 114.830 (22.726) ***
Afternoon peak -19.248 (1.474) *** 32.898 (2.300) *** 350.333 (21.615) ***
Afternoon valley -17.353 (1.465) *** 26.432 (2.291) *** 286.268 (21.268) ***
Night -190.052 (1.954) *** -404.928 (3.278) *** -1690.551 (30.204) ***
Night valley -128.461 (1.766) *** -100.212 (2.528) *** -214.808 (21.662) ***
Year 2019 1.846 (0.827) ** -7.982 (1.308) *** 12.130 (11.154)
Friday 4.843 (1.201) *** 7.624 (1.862) *** 27.000 (16.188)
Saturday -52.273 (1.281) *** -81.128 (1.973) *** -326.066 (17.207) ***
Sunday -97.262 (1.414) *** -154.568 (2.182) *** -652.903 (19.453) ***
Scooter trips destinations 45.404 (12.498) *** 99.799 (4.007) *** -136.663 (47.625) ***
Metro 14.584 (0.904) *** 2.883 (1.329) ** 119.174 (20.697) ***
Log(working population) 17.410 (0.507) *** 47.739 (1.309) *** 183.321 (10.685) ***
Relative educational level 18.830 (0.566) *** 36.069 (0.852) *** 344.369 (9.459) ***
*p $<$ 0.10; **p $<$ 0.05; ***p $<$ 0.01. Z-values are shown in parenthesis.
Table 9: Marginal effects of negative binomial models for evaluating generation of bus and metro trips
@@
Morning peak 0.553 (45.315) *** 0.229 (19.929) *** 0.003 (0.087)
Lunch -0.160 (-11.900) *** 0.018 (1.401) 0.166 (5.277) ***
Afternoon peak -0.142 (-11.869) *** 0.173 (15.519) *** 0.487 (17.079) ***
Afternoon valley -0.118 (-10.011) *** 0.173 (15.585) *** 0.438 (15.381) ***
Night -1.510 (-116.152) *** -1.915 (-162.572) *** -1.955 (-71.144) ***
Night valley -1.144 (-88.904) *** -0.584 (-48.888) *** -0.434 (-14.533) ***
@@
Friday 0.034 (3.518) *** 0.030 (3.337) *** 0.036 (1.621)
Saturday -0.478 (-47.719) *** -0.481 (-51.589) *** -0.475 (-20.657) ***
Sunday -0.879 (-83.864) *** -0.898 (-93.109) *** -0.937 (-39.569) ***
Treatment period
Year 2019 -0.022 (-3.376) *** -0.056 (-8.641) *** -0.206 (-13.259) ***
Treatment zones
Scooter trips destination 0.866 (8.604) *** 0.229 (12.127) *** 0.781 (11.980) ***
@@
Log(working population) 0.273 (70.063) *** 0.379 (72.816) *** 0.170 (13.165) ***
Relative educational level 0.105 (25.776) *** 0.145 (37.522) *** 0.313 (28.425) ***
@@
Table 10: Results of negative binomial models for evaluating generation of bus trips
@@
Morning peak 46.4909 (1.073)*** 20.1829 (1.019)*** 0.4706 (5.425)
Lunch -13.4742 (1.135)*** 1.5496 (1.106) 30.8586 (5.866)***
Afternoon peak -11.9116 (1.006)*** 15.2159 (0.984)*** 90.5389 (5.473)***
Afternoon valley -9.9645 (0.997)*** 15.2425 (0.982)*** 81.4608 (5.440)***
Night -127.0872 (1.289)*** -168.4752 (1.298)*** -363.5232 (6.644)***
Night valley -96.2803 (1.201)*** -51.3411 (1.084)*** -80.6726 (5.657)***
Year 2019 -1.8997 (0.563)*** -4.8398 (0.561)*** -38.3488 (2.944)***
Friday 2.8771 (0.818)*** 2.6661 (0.799)** 6.7609 (4.173)
Saturday -40.2679 (0.874)*** -42.2632 (0.848)*** -88.2394 (4.411)***
Sunday -73.7713 (0.966)*** -78.9807 (0.936)*** -174.2467 (4.877)***
Scooter trips destinations 72.8203 (8.481)*** 20.1006 (1.663)*** 145.2156 (12.422)***
Metro 12.0656 (0.610)*** -8.0391 (0.582)*** 91.3796 (5.520)***
log(Working population) 23.0446 (0.350)*** 33.3239 (0.492)*** 31.7262 (2.466)***
Relative educational level 8.8441 (0.347)*** 12.7893 (0.349)*** 58.2051 (2.209)***
@@
Table 11: Marginal effects of negative binomial models for evaluating generation of bus trips
@@
Morning peak 0.350 (11.775) *** -0.114 (-2.537) **
Lunch -0.100 (-3.164) *** 0.202 (4.242) ***
Afternoon peak 0.323 (11.080) *** 0.672 (15.188) ***
Afternoon valley 0.216 (7.450) *** 0.535 (12.125) ***
Night -2.831 (-79.758) *** -2.950 (-56.509) ***
Night valley -0.566 (-18.605) *** -0.306 (-6.834) ***
@@
Friday 0.066 (2.758) *** 0.042 (1.159)
Saturday -0.478 (-18.935) *** -0.653 (-17.247) ***
Sunday -0.899 (-33.717) *** -1.305 (-33.075) ***
Treatment period
Year 2019 -0.091 (-5.187) *** -0.043 (-1.696)
Treatment zones
Scooter trips destination 0.714 (17.166) *** 0.385 (1.456)
Infrastructure
Metro

### v2

@@ 6 Discussion
Our results suggest both dynamics can coexist within the same city but in different zones: substitution dominates where abundant transit offers redundant options, while complementarity emerges where e-scooters address genuine accessibility gaps. These findings underscore the importance of tailoring micromobility policy to local transit conditions rather than assuming uniform impacts across urban areas.
Our findings reveal that shared e-scooter impacts on public transport are fundamentally shaped by urban context, transport mode, and local accessibility conditions. Rather than uniform substitution or complementarity, we observe spatially differentiated patterns that account for the varied findings in existing literature.
Spatial Heterogeneity Explains Mixed Literature Findings
The coexistence of substitution effects in central areas (-23.87% in total boardings) and complementarity in intermediate zones (+33.64% in boardings) within the same urban system resolves the apparent contradiction in previous studies. Research reporting predominantly substitutive effects (Lu et al., 2024; Laa and Leth, 2020) may have focused on high-density areas similar to our Central Region, while studies emphasizing complementarity (Bai and Jiao, 2020; Vinagre Díaz et al., 2023) may have examined areas resembling our Intermediate Region. However, the significant combined bus-metro effect without corresponding individual mode significance in central areas suggests caution in interpreting underlying mechanisms.
This spatial variation aligns with findings from Oslo (Aarhaug et al., 2023) and Indianapolis (Luo et al., 2021), where both relationships occurred simultaneously within the same city. Our contribution lies in systematically demonstrating how built environment characteristics (transit density, trip distances, and accessibility levels) predict which relationship dominates in different urban areas.
Mode-Specific Effects Reflect Service Characteristics
Metro services show stronger complementary relationships with e-scooters than bus services across all regions. The significant increase in metro boardings in the Intermediate Region (+9.77%) contrasts with neutral bus effects, suggesting that fixed-route rail services benefit more from first/last-mile connections than flexible bus networks. This pattern aligns with Berlin findings where e-scooter demand increased near new metro stations (Weschke, 2023).
Bus services’ consistent neutrality across regions may reflect Santiago’s extensive, high-frequency network characteristics that make them less susceptible to e-scooter substitution. The comprehensive coverage and frequent stops of integrated bus systems may provide sufficient flexibility to maintain ridership even as alternative short-distance options emerge.
Urban Structure Drives Access Patterns
The strongest effects occur in the Intermediate Region, characterized by moderate transit density and intermediate trip distances. These conditions create optimal environments for e-scooters to bridge accessibility gaps without competing directly with efficient transit services. Central areas show substitution effects in combined bus and metro trips (-23.87%), though individual mode effects are not statistically significant, potentially reflecting statistical power limitations when disaggregating data or effects occurring primarily in multimodal trips. Peripheral areas show no significant effects, likely because trip distances exceed practical e-scooter range.
@@
These patterns suggest spatially differentiated strategies rather than uniform policies. Central areas could benefit from dynamic pricing that encourages e-scooter use during peak congestion periods while protecting transit during off-peak hours. Intermediate areas should focus on integration through coordinated infrastructure and payment systems to strengthen the observed complementary effects. Peripheral areas may require targeted deployment strategies to address mobility gaps.
The apparent substitution effects observed in central areas could potentially benefit system efficiency if they reduce peak-hour transit congestion, though this interpretation requires further investigation given the statistical patterns observed. Rather than restricting e-scooters, policies might consider strategic approaches to optimize mobility systems, while recognizing the need for additional research to understand underlying mechanisms.
Limitations and Future Research
Our analysis operates at the zonal level due to data integration constraints, potentially masking individual-level heterogeneity. The focus on Santiago limits generalizability, though the systematic spatial approach provides a framework for examining other contexts. Short-term effects may not capture longer-term adaptation patterns as users optimize multimodal journeys.
Future research should examine temporal variations to understand whether substitution effects help manage peak-hour congestion or create capacity challenges. Integration of individual-level data across transport modes would enable more precise understanding of modal shift dynamics and support development of adaptive policies that respond to real-time demand patterns across different urban contexts.
@@ 7 Conclusions
This study demonstrates spatially heterogeneous effects of shared e-scooter introduction on public transport usage in Santiago, Chile. Central areas experienced substitution effects (23.87% reduction in total boardings), intermediate areas showed complementary effects with increases in trip generation (33.64%) and attraction (4.08%), while peripheral areas exhibited no significant impacts.
These spatially differentiated patterns help explain the divergent findings in existing literature and underscore the need for context-sensitive policies that recognize e-scooters’ variable roles across urban environments. Metro services demonstrated stronger complementary relationships with e-scooters than bus services, particularly for first/last-mile connections. Our automated clustering methodology provides a replicable framework for future quasi-experimental studies, contributing to more rigorous comparative research across different urban contexts.
@@ Appendix C Results of Negative Binomial Models for Evaluating Generation of Public Transport Trips
Morning peak 0.5620 (25.938) *** 0.2413 (10.778) *** -0.0879 (-1.548)
Lunch -0.5106 (-23.628) *** -0.3235 (-14.510) *** -0.1018 (-1.798) *
Afternoon peak -0.0617 (-2.848) ** 0.3198 (14.506) *** 0.5771 (10.277) ***
Afternoon valley 0.0017 (0.079) 0.2867 (13.044) *** 0.4843 (8.656) ***
Night -1.8128 (-81.101) *** -2.3724 (-105.902) *** -2.3969 (-41.747) ***
Night valley -1.3019 (-59.505) *** -0.6664 (-29.799) *** -0.4283 (-7.574) ***
@@
Friday 0.0714 (4.163) *** 0.0561 (3.223) ** 0.0550 (1.233)
Saturday -0.4290 (-24.587) *** -0.4753 (-26.956) *** -0.5434 (-12.099) ***
Sunday -0.8734 (-48.838) *** -0.9535 (-52.715) *** -1.1174 (-24.561) ***
Difference-in-differences variables
Post 0.0117 (0.992) -0.0520 (-3.942) *** -0.0114 (-0.358)
Treatment_zones -0.3488 (-4.707) *** -0.2635 (-8.019) *** -0.4151 (-4.306) ***
Treatment_zones x post 0.1854 (1.873) * 0.3583 (11.194) *** -0.2491 (-2.265) **
@@
Log(working population) 0.1977 (28.755) *** 0.3562 (38.206) *** 0.3524 (13.045) ***
Relative educational level 0.2488 (29.738) *** 0.3741 (36.668) *** 0.6883 (30.801) ***
@@
Table 10: Results of negative binomial models for evaluating generation of bus and metro trips.
@@
Time periods
Morning peak 158.9475 (24.731) *** 120.5449 (10.701) *** -164.4222 (-1.547)
Lunch -144.4176 (-22.934) *** -161.6223 (-14.338) *** -190.3694 (-1.796) *
Afternoon peak -17.4446 (-2.846) ** 159.7837 (14.330) *** 1079.1796 (9.799) ***
Afternoon valley 0.4828 (0.079) 143.2316 (12.897) *** 905.5947 (8.365) ***
Night -512.7257 (-64.484) *** -1185.1764 (-76.167) *** -4482.0023 (-30.069) ***
Night valley -368.2161 (-51.614) *** -332.9000 (-28.389) *** -800.8503 (-7.406) ***
Type of day
Friday 20.1870 (4.157) *** 28.0295 (3.220) ** 102.7591 (1.232)
Saturday -121.3347 (-23.900) *** -237.4285 (-26.039) *** -1016.1518 (-11.555) ***
Sunday -247.0365 (-44.296) *** -476.3533 (-46.916) *** -2089.4423 (-20.936) ***
Difference-in-differences variables
Post 3.2993 (0.992) -25.9553 (-3.938) *** -21.2911 (-0.357)
Treatment zones -98.6635 (-4.700) *** -131.6356 (-7.976) *** -776.2562 (-4.282) ***
Treatment zones × post 52.4331 (1.873) * 178.9864 (11.069) *** -465.8094 (-2.262) **
Infrastructure
Metro 15.5300 (4.362) *** 88.7940 (13.095) *** 428.1174 (4.156) ***
Sociodemographic information
Log(working population) 55.9234 (27.747) *** 177.9331 (36.064) *** 658.8805 (12.011) ***
Relative educational level 70.3743 (28.174) *** 186.9059 (33.840) *** 1287.0477 (23.736) ***
*p $<$ 0.10; **p $<$ 0.05; ***p $<$ 0.01. Z-values are shown in parentheses.
Table 11: Marginal effects of negative binomial models for evaluating generation of metro and bus trips by region.
@@
Morning peak 0.5526 (45.303) *** 0.2293 (19.913) *** 0.0924 (2.081) **
Lunch -0.1522 (-11.824) *** 0.0150 (1.190) 0.1693 (5.386) ***
Afternoon peak -0.1416 (-11.863) *** 0.1739 (15.604) *** 0.4935 (17.313) ***
Afternoon valley -0.1180 (-9.971) *** 0.1737 (15.623) *** 0.4439 (15.586) ***
Night -1.5104 (-116.104) *** -1.9181 (-162.780) *** -1.9647 (-71.540) ***
Night valley -1.1450 (-88.931) *** -0.5830 (-48.824) *** -0.4331 (-14.516) ***
@@
Friday 0.0346 (3.561) *** 0.0307 (3.381) *** 0.0332 (1.482)
Saturday -0.4784 (-47.678) *** -0.4809 (-51.631) *** -0.4736 (-20.614) ***
Sunday -0.8771 (-83.860) *** -0.9001 (-93.325) *** -0.9397 (-39.669) ***
Difference-in-differences variables
Post -0.0204 (-3.025) ** -0.0387 (-5.752) *** -0.1795 (-11.306) ***
Treatment_zones 0.1019 (2.577) ** -0.1688 (-10.853) *** 0.4402 (8.800) ***
Treatment_zones x post 0.0511 (0.937) -0.0060 (-0.339) 0.1121 (1.689) *
@@
Log(working population) 0.2742 (70.098) *** 0.3806 (72.965) *** 0.2040 (14.851) ***
Relative educational level 0.1038 (24.288) *** 0.2075 (42.176) *** 0.2777 (23.306) ***
@@
Table 12: Results of negative binomial models for evaluating generation of bus trips.
@@
Time periods
Morning peak 46.5049 (43.318) *** 20.1732 (19.789) *** -0.4427 (-0.081)
Lunch -13.3951 (-11.796) *** 1.3160 (1.190) 31.6000 (5.368) ***
Afternoon peak -11.9123 (-11.831) *** 15.3047 (15.546) *** 92.0866 (16.749) ***
Afternoon valley -9.9317 (-9.953) *** 15.2853 (15.560) *** 82.8260 (15.162) ***
Night -127.1050 (-98.592) *** -168.7648 (-129.826) *** -366.6060 (-54.728) ***
Night valley -96.3604 (-80.193) *** -51.2927 (-47.290) *** -80.8218 (-14.244) ***
Type of day
Friday 2.9136 (3.560) *** 2.7022 (3.380) *** 6.2007 (1.482)
Saturday -40.2563 (-46.023) *** -42.3135 (-49.867) *** -88.3664 (-19.964) ***
Sunday -73.8129 (-76.349) *** -79.1927 (-84.575) *** -175.3454 (-35.764) ***
Difference-in-differences variables
Post -1.7151 (-3.024) ** -3.4088 (-5.749) *** -33.5015 (-11.165) ***
Treatment zones 8.5729 (2.577) ** -14.8500 (-10.827) *** 82.1445 (8.674) ***
Treatment zones x post 4.2967 (0.937) -0.5301 (-0.339) 20.9254 (1.688) *
Infrastructure
Metro 12.0582 (19.742) *** -4.8133 (-7.972) *** 89.7584 (16.228) ***
Sociodemographic information
Log(working population) 23.0765 (65.909) *** 33.4870 (67.721) *** 38.0689 (14.385) ***
Relative educational level 8.7374 (24.047) *** 18.2544 (40.908) *** 51.8139 (22.186) ***
*p $<$ 0.10; **p $<$ 0.05; ***p $<$ 0.01. Z-values are shown in parentheses.
Table 13: Marginal effects of negative binomial models for evaluating generation of bus trips by region.
Parameters Intermediate Region Central Region
Intercept 8.5065 (57.937) *** 6.2008 (26.294) ***
Time periods
Morning peak 0.2523 (6.568) *** -0.1863 (-3.845) ***
Lunch -0.3441 (-8.987) *** -0.0803 (-1.607)
Afternoon peak 0.4737 (12.547) *** 0.6747 (13.975) ***
Afternoon valley 0.3013 (8.039) *** 0.5419 (11.254) ***
Night -3.4498 (-90.987) *** -3.4877 (-70.918) ***
Night valley -0.5930 (-15.565) *** -0.3613 (-7.497) ***
Type of day
Friday 0.0563 (2.253) ** 0.0552 (1.448)
Saturday -0.6095 (-20.330) *** -0.7370 (-19.158) ***
Sunday -1.1318 (-35.975) *** -1.4706 (-37.483) ***
Difference-in-differences variables
Post 0.0274 (1.161) -0.0488 (-1.840) *
Treatment_zones 0.4337 (7.949) *** -0.2887 (-3.155) **
Treatment_zones x post 0.1012 (1.996) ** 0.1024 (0.811)
Sociodemographic information
Log(working population) -0.1169 (-6.742) *** 0.2159 (7.932) ***
Relative educational level 0.1901 (10.186) *** 0.4170 (20.720) ***
Model performance
Nº observations 6,635 2,242
Log-likelihood -5.3314e+04 -1.9238e+04
@@
Table 14: Results of negative binomial models for evaluating generation of metro trips.
@@
Morning peak 524.6563 (6.521) *** -820.6119 (-3.823) ***
Lunch -715.6997 (-8.867) *** -353.5639 (-1.666) *
Afternoon peak 985.0912 (12.214) *** 2971.7075 (13.117) ***
Afternoon valley 626.5734 (7.945) *** 2386.4923 (10.795) ***
Night -7174.3981 (-54.013) *** -15360.0000 (-41.987) ***
Night valley -1233.1786 (-14.965) *** -1591.3766 (-7.378) ***
@@
Friday 137.8575 (2.250) ** 243.0036 (1.446)
Saturday -1267.5468 (-19.168) *** -3246.0942 (-17.468) ***
Sunday -2353.7217 (-30.772) *** -6476.8729 (-28.807) ***
Difference-in-differences variables
Post 56.8927 (1.161) -215.1262 (-1.837) *
Treatment destination zones 901.9012 (7.843) *** -1271.6845 (-3.142) ***
Treatment zones × post 210.5348 (1.994) ** 450.9161 (0.811)